# A Survey of Efficient Reasoning for Large Reasoning Models: Language, Multimodality, and Beyond

### v1

@@ 6.2.1 Linear Sequence Modeling
Linear attention methods Qin et al. (2024c) exploit the “right-product kernel trick”, wherein the initial computation of key-value products circumvents the quadratic cost typically incurred by query-key interactions. For example, vanilla linear attention (Katharopoulos et al., 2020) replaces traditional attention (Vaswani et al., 2017) with kernel-based approximations, thereby achieving linear computational complexity (Shen et al., 2024). Various enhancements have been developed to further boost efficiency. TransNormerLLM (Qin et al., 2023) introduces Lightning Attention (Qin et al., 2024e), which optimizes I/O operations to expedite processing (Qin et al., 2024b), while Lightning Attention-2 (Qin et al., 2024d) refines block-wise computations for superior performance in autoregressive settings. Moreover, sequence parallelism has been investigated to extend the capability of linear attention models for handling long sequences across extensive clusters. LASP (Sun et al., 2024c) was the first to integrate sequence parallelism into these methods, and its successor LASP-2 (Sun et al., 2025b) further refines the approach by reorganizing both computational and communication workflows. Additionally, Minimax-01 (Li et al., 2025a) adapts the Lightning Attention and LASP-series strategies to a massive MoE language model with 456 billion parameters, highlighting its potential for commercial deployment.
@@
Other innovations in linear attention mechanisms include RetNet (Sun et al., 2023), which introduces a retention mechanism that supports parallel training without sacrificing linear-time inference. Gated Linear Attention (GLA) (Yang et al., 2023) leverages a data-independent gating scheme to enhance the sequence modeling ability and hardware efficiency, while Gated Slot Attention (GSA) (Zhang et al., 2024c) employs a bounded-memory slot control strategy to improve recall in tasks with extended contexts. Furthermore, approaches such as Test-Time Training (TTT) (Sun et al., 2024e), Titans (Behrouz et al., 2024), and Gated-DeltaNet (Yang et al., 2024e, d) propose update rules that allow models to adapt dynamically during inference. Despite their differing gating and updating strategies, these methods generally depend on a fixed-size memory state. In contrast, MoM (Du et al., 2025) expands the RNN memory state using “sparse memory” with multiple memory units managed by a router module.
@@
State Space Models (SSMs) (Gu et al., 2022b, a; Gupta et al., 2022; Gu and Dao, 2023) represent a promising approach for efficient sequence modeling. The latest variant, Mamba2 (Dao and Gu, 2024), integrates a linear attention-like mechanism to enhance hardware efficiency, utilizing state-space duality to support parallel computation while maintaining recurrent inference capabilities. In addition to SSMs, linear RNN-based methods such as RWKV (Peng et al., 2024, 2025a), HGRN (Qin et al., 2024g), and its successor HGRN2 (Qin et al., 2024f) have also demonstrated effectiveness in improving the efficiency of sequence modeling.
@@ 6.2.2 Sparse Attention
Sparse attention presents another effective approach for managing long sequences and mitigating the quadratic complexity of self-attention mechanisms, while still maintaining model performance. Sliding window attention limits queries to a predetermined local context. Approaches such as StreamingLLM (Xiao et al., 2023b), MoA (Fu et al., 2024a; Qu et al., 2024b), and DuoAttention (Xiao et al., 2024) integrate local attention with attention sinks to facilitate efficient long-sequence processing. Longformer (Beltagy et al., 2020) alternates between local attention and global tokens to improve contextual understanding.
NSA (Yuan et al., 2025a) adopts a dynamic hierarchical sparse strategy, combining coarse-grained token compression with fine-grained token selection, ensuring both global context retention and local accuracy. MoBA (Lu et al., 2025a) divides the context into blocks and uses a dynamic gating mechanism to route query tokens to the most relevant KV blocks.
@@ 6.3 Linearization
The linearization of large language models converts pre-trained standard models into linear recurrent structures, enhancing deployment efficiency. Liger (Lan et al., 2025) modifies pre-trained LLMs into gated linear recurrent models by adapting key matrix weights, removing the need for extra parameters. LoLCATs Zhang et al. (2024a) advances LLM linearization by replacing softmax attention with trained linear approximations and enhancing model quality using LoRA. Llamba (Bick et al., 2025) distills Llama-3.x models into the Mamba architecture, achieving high inference throughput and efficiency with minimal training data through MOHAWK (Bick et al., 2024). LightTransfer (Zhang et al., 2025b) decreases KV-cache memory demands in long-context LLMs by substituting certain attention layers with streaming attention. MOHAWK (Bick et al., 2024) facilitates the distillation of pre-trained Transformers into subquadratic models like SSMs through a structured three-phase process. Multimodal Mamba (Liao et al., 2025) builds linear-complexity multimodal state space models from existing MLLMs via progressive distillation, lowering computational costs.
@@ 6.4 Efficient Reasoning with Subquadratic Attention
Think Slow Fast (Paliotta et al., 2025) investigate subquadratic models for efficient reasoning under fixed compute budgets, demonstrating that distilling Mamba models from Transformers enables faster multi-path CoT generation and improved performance on MATH and GSM8K. Their findings highlight the potential of architectural innovations to enhance reasoning efficiency in resource-constrained environments. CRQs (Yehudai et al., 2025) examine the expressiveness of Transformers, RNNs, and CoT-augmented models on Compositional Reasoning Questions (CRQs), revealing inherent trade-offs: Transformers require logarithmic depth, RNNs depend on logarithmic embedding dimensions, and CoT-augmented Transformers scale with linearly many tokens. This work provides a formal framework for comparing model capabilities in multi-step reasoning tasks. Cosmos-Reason1 (Azzolini et al., 2025) proposes to use the hybrid Mamba-MLP-Transformer backbone architecture to enable efficient Physical AI reasoning (with physical common sense, embodied reasoning, and logical inference). The hybrid architecture was demonstrated to be both efficient and well-performed on long CoT reasoning.
@@ 6.5 Summary and Outlook
Looking ahead, the integration of subquadratic attention with emerging reasoning paradigms, such as hybrid architectures and adaptive retrieval mechanisms, holds great potential for enhancing the efficiency and scalability of large-scale reasoning models. Additionally, novel techniques in large language models, including block diffusion (Arriola et al., 2025) and large language diffusion models (Arriola et al., 2025; Nie et al., 2025), merit further exploration. These advancements may introduce new opportunities while also presenting unique challenges in the pursuit of more efficient and capable reasoning systems.
@@ 7.1 Efficient Multimodal Reasoning and Video Reasoning
Recently, researchers Meng et al. (2025); Liu et al. (2025d); Huang et al. (2025c); Pan et al. (2025); Zhou et al. (2025a); Peng et al. (2025c); Thawakar et al. (2025); Liu et al. (2025a); Yao et al. (2024a); Peng et al. (2025b) have demonstrated that o1-like CoT reasoning also plays a significant role in multimodal reasoning Liu et al. (2023a); Liang et al. (2024). In addition, Zhao et al. (2025) further extend the reasoning ability to emotion recognition in the video. In contrast to directing output in the traditional multimodal large language models Wang et al. (2024a); Chen et al. (2024b); Team (2023); Zhang et al. (2023a), the augmented reasoning capabilities for multimodal large reasoning models enable a nuanced dissection of modal contributions. However, efficient reasoning is under-explored in multimodal reasoning and video reasoning. The complex image and video information contains more noisy information than language which limits efficient reasoning. Moreover, reasoning models in these domains remain susceptible to overthinking and excessive computational costs Wang et al. (2025c). As shown in Figure 9, inefficient reasoning often leads to unnecessary complexity and convoluted paths to the correct answer.
@@
To solve this problem, the current works all follow two key principles: 1) Different problem types may demand distinct forms of reasoning capabilities; 2) The complexity of the reasoning process should align with the inherent difficulty of the problem. Self-structured Chain of Thought (SCoT) framework Xiang et al. (2025) has taken a significant step in addressing these issues by decomposing reasoning tasks into atomic, semantically meaningful steps. This approach ensures that the reasoning process is not only efficient but also adaptable to the complexity of the task at hand, particularly in multimodal settings Xia et al. (2024). For instance, simpler tasks like image captioning may rely on fewer atomic steps Stefanini et al. (2021), while more intricate tasks, such as emotion recognition in video, may demand a deeper understanding of temporal and visual features, requiring a more sophisticated chain of reasoning Canal et al. (2022). The Adaptive-Length Chain-of-Thought Distillation (AL-CoTD) framework Peng et al. (2025b) further refines this process by dynamically adjusting the length of reasoning chains according to task complexity. This addresses the issue of overthinking, which is particularly prevalent in multimodal and video reasoning tasks.
@@
Traditional multimodal large language models often treat reasoning as a uniform process, regardless of the input type. In contrast, integrating dynamic reasoning strategies across different modalities (e.g., textual, image, and video data) allows for more efficient problem-solving by adjusting the cognitive complexity to match the nature of the inputs. This approach promises to handle the noise and complexity inherent in different modalities of data while maintaining computational efficiency. By balancing expressiveness with efficiency, future multimodal reasoning models can adapt to various tasks, using simpler reasoning for less complex problems and more sophisticated methods for challenging ones, ultimately driving more powerful and efficient multimodal solutions Liang et al. (2024); Meng et al. (2025); Liu et al. (2025d).
@@ 7.2 Efficient Test-time Scaling and Infinity Thinking
Test-time scaling is a direct method for extending thinking time and improving response quality and model performance in LRMs and LLMs Snell et al. (2024); Bi et al. (2024); Liu et al. (2025b), and is typically classified as parallel sampling and sequential revision. For parallel sampling, the scaling can be conducted by extending the search width, such as Best-of-N sampling Cobbe et al. (2021b), Self-Consistency Wang et al. (2022a), and minimum Bayes risk decoding Wu et al. (2024); Heineman et al. (2024). However, these approaches necessitate a fixed number of samples per query, irrespective of complexity. This leads to computational waste on simpler queries and potentially insufficient exploration for complex ones. This kind of inefficient reasoning can be mitigated by developing confidence-based methods to address queries of varying difficulty Huang et al. (2025a) or adaptive sampling strategies Wan et al. (2024); Wang et al. (2025d); Li et al. (2024b).
For sequential revision, scalability can be achieved by extending the search depth for reasoning, incorporating methods such as debating Liang et al. (2023); Du et al. (2023), self-correction Lin et al. (2024b), and self-critique Yu et al. (2024b); Su et al. (2024d). These techniques allow the reasoning process to expand into increasingly longer sequences. In extreme cases, this reasoning process may even become infinite. Yan et al. (2025) have addressed this emerging challenge for large retrieval models (LRMs) by transforming reasoning into an iterative process with intermediate summarizations. This approach interleaves short reasoning segments with concise progress summaries, allowing the depth of reasoning to be significantly extended while maintaining manageable computational costs. Similarly, Yang et al. (2025a) propose a reduction rule that reduces context length during the standard iterative next-token generation, applying it whenever feasible to optimize the process.
@@ 7.3 Efficient and Trustworthy Reasoning
Emerging large reasoning models, e.g., OpenAI o1 and DeepSeek-R1, generate long and structured CoT steps, resulting in astonishing performance Su et al. (2024b); Li et al. (2025b). However, such long CoT steps bring new challenges to the trustworthiness of LRMs, including safety and reliability Liu et al. (2023b); Yao et al. (2024b).
For the safety concern Ren et al. (2024); Qian et al. (2024a); Hu et al. (2024b), researchers Zhou et al. (2025b); Jiang et al. (2025) discover that the safe rate of the LRMs’ thinking process is lower than that of the final answers. When presented with harmful queries, LRMs engage in reasoning and generate relevant content, potentially exposing sensitive information even if the final response is ultimately safe. To mitigate this, deliberative alignment Guan et al. (2024) and X-Boundary Lu et al. (2025b) offer distinct approaches to enhance LRM safety. Exploring efficient and inherently safe reasoning mechanisms presents a compelling future direction, echoing the adage “he that talks much errs much”. For instance, latent space reasoning Deng et al. (2024); Su et al. (2024a); Sui et al. (2025); Kong et al. (2025) offers efficiency, while representation engineering Zou et al. (2023, 2024); Qian et al. (2024b); Chen et al. (2025a); Liu et al. (2024b) can ensure robust safety performance.
In terms of reliability, LLMs are known to suffer from both factuality and faithfulness hallucinations Huang et al. (2025b) across various applications. This problem is potentially amplified in LRMs, whose extended reasoning chains are inherently more susceptible to the accumulation of noisy and untrustworthy information. The increased complexity and length of these reasoning processes provide greater opportunity for errors and deviations from ground truth to accumulate. Furthermore, CoTs do not accurately reflect the model thinking process and bring additional uncertainty Lanham et al. (2023); Su et al. (2023); Turpin et al. (2023); Tanneru et al. (2024); Agarwal et al. (2024); Su et al. (2024c). These issues create a compounding effect, worsening the hallucination problem in both language-based and multi-modal contexts Qu et al. (2024c). Therefore, the development of efficient and reliable reasoning methodologies represents a particularly pressing research need in the age of LRMs.

### v2

@@ 6.2.1 Linear Sequence Modeling
Linear attention methods qin2024unlocking exploit the “right-product kernel trick”, wherein the initial computation of key-value products circumvents the quadratic cost typically incurred by query-key interactions. For example, vanilla linear attention (katharopoulos2020transformers) replaces traditional attention (vaswani2017attention) with kernel-based approximations, thereby achieving linear computational complexity (shen2024scaling). Various enhancements have been developed to further boost efficiency. TransNormerLLM (qin2023scaling) introduces Lightning Attention (qin2024various), which optimizes I/O operations to expedite processing (qin2024transnormerllm), while Lightning Attention-2 (qin2024lightning) refines block-wise computations for superior performance in autoregressive settings. Moreover, sequence parallelism has been investigated to extend the capability of linear attention models for handling long sequences across extensive clusters. LASP (sun2024linear) was the first to integrate sequence parallelism into these methods, and its successor LASP-2 (sun2025lasp) further refines the approach by reorganizing both computational and communication workflows. Additionally, Minimax-01 (li2025minimax) adapts the Lightning Attention and LASP-series strategies to a massive MoE language model with 456 billion parameters, highlighting its potential for commercial deployment.
@@
Other innovations in linear attention mechanisms include RetNet (sun2023retentive), which introduces a retention mechanism that supports parallel training without sacrificing linear-time inference. Gated Linear Attention (GLA) (yang2023gated) leverages a data-independent gating scheme to enhance the sequence modeling ability and hardware efficiency, while Gated Slot Attention (GSA) (zhang2024gsa) employs a bounded-memory slot control strategy to improve recall in tasks with extended contexts. Furthermore, approaches such as Test-Time Training (TTT) (sun2024learning), Titans (behrouz2024titans), and Gated-DeltaNet (yang2024parallelizing; yang2024gated) propose update rules that allow models to adapt dynamically during inference. Despite their differing gating and updating strategies, these methods generally depend on a fixed-size memory state. In contrast, MoM (du2025mom) expands the RNN memory state using “sparse memory” with multiple memory units managed by a router module.
@@
State Space Models (SSMs) (gu2022efficiently; gu2022parameterization; gupta2022diagonal; gu2023mamba) represent a promising approach for efficient sequence modeling. The latest variant, Mamba2 (dao2024transformers), integrates a linear attention-like mechanism to enhance hardware efficiency, utilizing state-space duality to support parallel computation while maintaining recurrent inference capabilities. In addition to SSMs, linear RNN-based methods such as RWKV (peng2024eagle; peng2025rwkv7gooseexpressivedynamic), HGRN (qin2024hierarchically), and its successor HGRN2 (qin2024hgrn2) have also demonstrated effectiveness in improving the efficiency of sequence modeling.
@@ 6.2.2 Sparse Attention
Sparse attention presents another effective approach for managing long sequences and mitigating the quadratic complexity of self-attention mechanisms, while still maintaining model performance. Sliding window attention limits queries to a predetermined local context. Approaches such as StreamingLLM (streaming), MoA (moa; qu2024llama), and DuoAttention (xiao2024duoattention) integrate local attention with attention sinks to facilitate efficient long-sequence processing. Longformer (beltagy2020longformer) alternates between local attention and global tokens to improve contextual understanding.
NSA (yuan2025native) adopts a dynamic hierarchical sparse strategy, combining coarse-grained token compression with fine-grained token selection, ensuring both global context retention and local accuracy. MoBA (lu2025moba) divides the context into blocks and uses a dynamic gating mechanism to route query tokens to the most relevant KV blocks.
@@ 6.3 Linearization
The linearization of large language models converts pre-trained standard models into linear recurrent structures, enhancing deployment efficiency. Liger (lan2025liger) modifies pre-trained LLMs into gated linear recurrent models by adapting key matrix weights, removing the need for extra parameters. LoLCATs zhang2024lolcats advances LLM linearization by replacing softmax attention with trained linear approximations and enhancing model quality using LoRA. Llamba (bick2025llamba) distills Llama-3.x models into the Mamba architecture, achieving high inference throughput and efficiency with minimal training data through MOHAWK (bick2024transformers). LightTransfer (zhang2025lighttransferlongcontextllmsecretly) decreases KV-cache memory demands in long-context LLMs by substituting certain attention layers with streaming attention. MOHAWK (bick2024transformers) facilitates the distillation of pre-trained Transformers into subquadratic models like SSMs through a structured three-phase process. Multimodal Mamba (liao2025multimodalmambadecoderonlymultimodal) builds linear-complexity multimodal state space models from existing MLLMs via progressive distillation, lowering computational costs.
@@ 6.4 Efficient Reasoning with Subquadratic Attention
Think Slow Fast (paliotta2025thinking) investigate subquadratic models for efficient reasoning under fixed compute budgets, demonstrating that distilling Mamba models from Transformers enables faster multi-path CoT generation and improved performance on MATH and GSM8K. Their findings highlight the potential of architectural innovations to enhance reasoning efficiency in resource-constrained environments. CRQs (yehudai2025compositionalreasoningtransformersrnns) examine the expressiveness of Transformers, RNNs, and CoT-augmented models on Compositional Reasoning Questions (CRQs), revealing inherent trade-offs: Transformers require logarithmic depth, RNNs depend on logarithmic embedding dimensions, and CoT-augmented Transformers scale with linearly many tokens. This work provides a formal framework for comparing model capabilities in multi-step reasoning tasks. Cosmos-Reason1 (azzolini2025cosmos) proposes to use the hybrid Mamba-MLP-Transformer backbone architecture to enable efficient Physical AI reasoning (with physical common sense, embodied reasoning, and logical inference). The hybrid architecture was demonstrated to be both efficient and well-performed on long CoT reasoning.
@@ 6.5 Summary and Outlook
Looking ahead, the integration of subquadratic attention with emerging reasoning paradigms, such as hybrid architectures and adaptive retrieval mechanisms, holds great potential for enhancing the efficiency and scalability of large-scale reasoning models. Additionally, novel techniques in large language models, including block diffusion (arriola2025block) and large language diffusion models (arriola2025block; nie2025large), merit further exploration. These advancements may introduce new opportunities while also presenting unique challenges in the pursuit of more efficient and capable reasoning systems.
@@ 7.1 Efficient Multimodal Reasoning and Video Reasoning
Recently, researchers meng2025mm; liu2025visual; huang2025vision; pan2025medvlm; zhou2025r1; peng2025lmm; thawakar2025llamav; liu2025seg; yao2024mulberry; skywork2025r1v have demonstrated that o1-like CoT reasoning also plays a significant role in multimodal reasoning liu2023mathematical; liang2024survey. In addition, zhao2025r1omniexplainableomnimultimodalemotion further extend the reasoning ability to emotion recognition in the video. In contrast to directing output in the traditional multimodal large language models wang2024qwen2; chen2024internvl; team2023internlm; zhang2023internlm, the augmented reasoning capabilities for multimodal large reasoning models enable a nuanced dissection of modal contributions. However, efficient reasoning is under-explored in multimodal reasoning and video reasoning. The complex image and video information contains more noisy information than language which limits efficient reasoning. Moreover, reasoning models in these domains remain susceptible to overthinking and excessive computational costs wang2025multimodalchainofthoughtreasoningcomprehensive. As shown in Figure 9, inefficient reasoning often leads to unnecessary complexity and convoluted paths to the correct answer.
@@
To solve this problem, the current works all follow two key principles: 1) Different problem types may demand distinct forms of reasoning capabilities; 2) The complexity of the reasoning process should align with the inherent difficulty of the problem. Self-structured Chain of Thought (SCoT) framework xiang2025atomicstepdecompositionenhance has taken a significant step in addressing these issues by decomposing reasoning tasks into atomic, semantically meaningful steps. This approach ensures that the reasoning process is not only efficient but also adaptable to the complexity of the task at hand, particularly in multimodal settings xia2024rule. For instance, simpler tasks like image captioning may rely on fewer atomic steps stefanini2021tellsurveydeeplearningbased, while more intricate tasks, such as emotion recognition in video, may demand a deeper understanding of temporal and visual features, requiring a more sophisticated chain of reasoning canal2022survey. The Adaptive-Length Chain-of-Thought Distillation (AL-CoTD) framework skywork2025r1v further refines this process by dynamically adjusting the length of reasoning chains according to task complexity. This addresses the issue of overthinking, which is particularly prevalent in multimodal and video reasoning tasks.
@@
Traditional multimodal large language models often treat reasoning as a uniform process, regardless of the input type. In contrast, integrating dynamic reasoning strategies across different modalities (e.g., textual, image, and video data) allows for more efficient problem-solving by adjusting the cognitive complexity to match the nature of the inputs. This approach promises to handle the noise and complexity inherent in different modalities of data while maintaining computational efficiency. By balancing expressiveness with efficiency, future multimodal reasoning models can adapt to various tasks, using simpler reasoning for less complex problems and more sophisticated methods for challenging ones, ultimately driving more powerful and efficient multimodal solutions liang2024survey; meng2025mm; liu2025visual.
@@ 7.2 Efficient Test-time Scaling and Infinity Thinking
Test-time scaling is a direct method for extending thinking time and improving response quality and model performance in LRMs and LLMs snell2024scaling; bi2024forest; liu2025iterative, and is typically classified as parallel sampling and sequential revision. For parallel sampling, the scaling can be conducted by extending the search width, such as Best-of-N sampling cobbe2021training, Self-Consistency wang2022selfconsis, and minimum Bayes risk decoding wu2024better; heineman2024improving. However, these approaches necessitate a fixed number of samples per query, irrespective of complexity. This leads to computational waste on simpler queries and potentially insufficient exploration for complex ones. This kind of inefficient reasoning can be mitigated by developing confidence-based methods to address queries of varying difficulty huang2025efficienttesttimescalingselfcalibration or adaptive sampling strategies wan2024dynamic; wang2025sampling; li2024escape.
For sequential revision, scalability can be achieved by extending the search depth for reasoning, incorporating methods such as debating liang2023encouraging; du2023improving, self-correction lin2024criticbench, and self-critique yu2024self; su2024timo. These techniques allow the reasoning process to expand into increasingly longer sequences. In extreme cases, this reasoning process may even become infinite. yan2025inftythink have addressed this emerging challenge for large retrieval models (LRMs) by transforming reasoning into an iterative process with intermediate summarizations. This approach interleaves short reasoning segments with concise progress summaries, allowing the depth of reasoning to be significantly extended while maintaining manageable computational costs. Similarly, yang2025pencil propose a reduction rule that reduces context length during the standard iterative next-token generation, applying it whenever feasible to optimize the process.
@@ 7.3 Efficient and Trustworthy Reasoning
Emerging large reasoning models, e.g., OpenAI o1 and DeepSeek-R1, generate long and structured CoT steps, resulting in astonishing performance su2024living; li2025llms. However, such long CoT steps bring new challenges to the trustworthiness of LRMs, including safety and reliability liu2023trustworthy; yao2024survey.
For the safety concern actor_attack; qian2024dean; vlsbench, researchers zhou2025hidden; jiang2025safechain discover that the safe rate of the LRMs’ thinking process is lower than that of the final answers. When presented with harmful queries, LRMs engage in reasoning and generate relevant content, potentially exposing sensitive information even if the final response is ultimately safe. To mitigate this, deliberative alignment guan2024deliberative and X-Boundary lu2025x offer distinct approaches to enhance LRM safety. Exploring efficient and inherently safe reasoning mechanisms presents a compelling future direction, echoing the adage “he that talks much errs much”. For instance, latent space reasoning deng2024explicit; su2024dualformercontrollablefastslow; sui2025metareasonerdynamicguidanceoptimized; kong2025scalable offers efficiency, while representation engineering zou2023representation; circuit_breaker; qian2024towards; chen2025seer; liu2024survey can ensure robust safety performance.
In terms of reliability, LLMs are known to suffer from both factuality and faithfulness hallucinations huang2025survey across various applications. This problem is potentially amplified in LRMs, whose extended reasoning chains are inherently more susceptible to the accumulation of noisy and untrustworthy information. The increased complexity and length of these reasoning processes provide greater opportunity for errors and deviations from ground truth to accumulate. Furthermore, CoTs do not accurately reflect the model thinking process and bring additional uncertainty lanham2023measuring; su2023efficient; turpin2023language; tanneru2024hardness; agarwal2024faithfulness; su2024conflictbank. These issues create a compounding effect, worsening the hallucination problem in both language-based and multi-modal contexts qu2024look. Therefore, the development of efficient and reliable reasoning methodologies represents a particularly pressing research need in the age of LRMs.